the total electric force on charge 1 measured counter-clockwise from the +x axis, in degrees? DQuestion 10 10.5 pts Charge 1 of +4 micro-coulombs and a mass of 0.3 kg lies on a horizontal surface and charge 2 of -5 micro-coulombs is placed directly 0.46 meters above this charge on the surface. What is the normal force, in Newtons, acting on the charge on the surface?
Answer #1

Since charges are unlike ,the electrostatic force is attractive,pulling the charge on the surface upwards.

Given q_{1}=4mu C=4*10^{-6}C

(2

distance, d-0.46m

Use Formula ,kq142

F 9 104105 10/(0.46m)

F=9*109 * 4 * 10-6 * 5 * 10-6/0.462

F 0.18/0.46

F-0.851N

=====================

Normal Force =Weight of charge at surface(acting downwards) - Force acting upwards

Normal Force =mg - F

N 0.3kg * 9.81-0.85 IN

ANSWER: N=2.092N

=====================